OpenStudy (buttermequeasy):

Solve for x: |x − 2| + 10 = 12

OpenStudy (karatechopper):

x=0

OpenStudy (anonymous):

|x-2| = 2 x-2= +-2 x=0,4

OpenStudy (anonymous):

x-2 + 10 = 12 x = 12-10 + 2 = 4 or x-2 = -2 x = 0 0,4